OpenStudy (anonymous):

A sphere with a radius of 4.8 centimeters is carved out of a right cone with a base radius of 8 centimeters and a height of 15 centimeters. What is the approximate volume of the remaining portion of the cone in terms of pi

jimthompson5910 (jim_thompson5910):

hint: think of the cross section of the cone as a triangle and the cross section of the sphere as a circle try to fit the circle into the triangle the best you can (max out the area of the circle based on the fixed triangle)

jimthompson5910 (jim_thompson5910):

something like this http://mathcentral.uregina.ca/QQ/database/QQ.09.07/h/linnea1.1.gif

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

129.32πcm3 147.46πcm3 172.54πcm3 194.68πcm3 the choices

OpenStudy (anonymous):

pi not ?

jimthompson5910 (jim_thompson5910):

what is the volume of the cone? what is the volume of the sphere?

OpenStudy (anonymous):

168.89 for cone??

jimthompson5910 (jim_thompson5910):

you're using this formula right? V = (1/3)*pi*r^2*h this is for the volume of a cone

OpenStudy (anonymous):

1005.31

jimthompson5910 (jim_thompson5910):

leave it in terms of pi

OpenStudy (anonymous):

960pi

jimthompson5910 (jim_thompson5910):

you should have 320pi not 960pi check your work again

OpenStudy (anonymous):

Idid 1/3*8^2*15

OpenStudy (anonymous):

oh i forgot to divide by 3

jimthompson5910 (jim_thompson5910):

now use this formula to find the volume of the sphere V = (4/3)*pi*r^3 leave the volume in terms of pi

jimthompson5910 (jim_thompson5910):

then you will do (volume of cone) - (volume of sphere)

OpenStudy (anonymous):

542.06

jimthompson5910 (jim_thompson5910):

I'm not getting that for the volume of the sphere

OpenStudy (anonymous):

no thats what i got for vol. of cone- sphere

OpenStudy (anonymous):

for vol. of sphere i got 147.46

jimthompson5910 (jim_thompson5910):

so, (volume of cone) - (volume of sphere) = (320pi) - (147.456pi) = 172.544pi

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np